OpenStudy (anonymous):

what is the simplified form of 4z^2-16z+15/2z^2+z-15

OpenStudy (lalaly):

\[\frac{4z^2-16z+15}{2z^2+z-15}\] can u factorise these two polynomials?

OpenStudy (anonymous):

yes, i factored them to: (4z-15)(z-1)/(2z-15)(z+1)

OpenStudy (lalaly):

thats not right ... \[=\frac{(2z-5)(2z-3)}{(z+3)(2z-5)}\]

OpenStudy (anonymous):

but i thought the sum of the two numbers has to equal the middle number...

OpenStudy (anonymous):

Working together, it takes two computers 12 minutes to send out a company’s email. If it takes the slower computer 30 minutes to do the job on its own, how long will it take the faster computer to do the job on its own? Do not do any rounding.

OpenStudy (lalaly):

yeah thats only when the coeffecient of x^2 or z^2 =1

OpenStudy (anonymous):

what is it equal to here? and how would i find out?